OpenStudy (anonymous):

Which of the following equations can be used to find the length of in the triangle below? A. (BC)2 = 102 + 302 B. BC = 30 - 10 C. (BC)2 = 302 - 102 D. BC = 30 + 10

OpenStudy (anonymous):

hweres the triangle

OpenStudy (anonymous):

it dosent show but is ba=10 ac=30 bc=??

OpenStudy (anonymous):

Which of the following equations can be used to find the length of in the triangle below?  A. (BC)2 = 102 + 302 B. BC = 30 - 10 C. (BC)2 = 302 - 102 D. BC = 30 + 10